# Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series

### v2

@@ 6 Groundwater data application
The file contains the proofs of the propositions presented in Section 4 and some technical results. It also documents the scalar filter algorithm of Peña and Tsay, (2021) and complementary numerical results, including the simulation results for autocovariance estimation and for data generated from a compositional time series model. Finally, it includes the analysis of the groundwater data via autoregressive modeling.
Supplementary material to “Temporal Wasserstein Imputation: Versatile Missing Data Imputation for Time Series”
This supplementary material contains three sections. Section S1 collects the proofs of the propositions presented in main text as well as some technical results. Section S2 documents the scalar filter algorithm of Peña and Tsay, (2021) for completeness. Section S3 contains complementary numerical results, including the simulation results for autocovariance estimation and imputing data generated from a compositional time series model. It also includes the analysis of autoregressive modeling using the imputed groundwater data.
@@
Since $\ell$ is convex, ′′$\ell^{\prime\prime}\geq 0$. From (S1.1)–(S1.3), it is readily seen that the Hessian matrix $\mathbf{H}=\partial^{2}F_{0}/\partial\mathbf{w}^{2}$ is weakly diagonally dominant with nonnegative diagonal entries. It follows from the Gershgorin circle theorem that $\mathbf{H}$ is positive semi-definite. Thus $F_{0}$ is convex in $\mathbf{w}$. ∎
Before proving Proposition 4, we first state our assumptions and establish Lemma S6.
The underlying time series $\{\tilde{x}_{t}\}$ is stationary and ergodic.
$\mathbb{E}|\tilde{x}_{t}|^{k}<\infty$ for some $k\geq 1$.
These assumptions are satisfied by many linear and nonlinear time series models.
@@
=$\mathcal{E}=\{\hat{\varphi}_{n}(\mathbf{q})\rightarrow\varphi(\mathbf{q})\mbox{ for all }\mathbf{q}\in\mathbb{Q}^{p}\}\cap\left\{\frac{1}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|\rightarrow\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|\right\},$
noting that $\mathbb{E}\|\mathbf{v}_{p-1}(\tilde{\mathbf{x}})\|<\infty$ by (A2). By the Ergodic theorem and (S1.6), $\mathbb{P}(\mathcal{E})=1$. Fix $\omega\in\mathcal{E}$ and let $\epsilon>0$ and $\mathbf{h}\in\mathbb{R}^{p}$ be arbitrary. Since $\varphi$ is uniformly continuous, there exists $\delta>0$ such that
@@
$|\hat{\varphi}_{n}(\mathbf{h})-\hat{\varphi}_{n}(\mathbf{q})|\leq\frac{\delta}{n-p+1}\sum_{t=p-1}^{n-1}\|\mathbf{v}_{t}(\tilde{\mathbf{x}})\|,$ (S1.8)
@@
where $\hat{\mathbf{w}}^{(i)}=(\hat{w}_{0}^{(i)},\ldots,\hat{w}_{n-1}^{(i)})$ and $\gamma_{\mathbf{x}}(h)=\frac{1}{1000}\sum_{i=1}^{1000}\gamma(\mathbf{x}^{(i)},h)$ with $\mathbf{x}^{(i)}=(x_{0}^{(i)},\ldots,x_{n-1}^{(i)})$.
Table S1 reports $\mathrm{RMSE}_{\gamma}(h)$ for $h=0,1,2$. For Model 1 (AR), most methods perform reasonably well. Linear interpolation produces accurate ACF estimates since the positive AR coefficient intrinsically favors linearity, while the scalar filter exhibits largest estimation errors. For Model 2 (ARMA), TWI improves from linear and Kalman initializations and yields more accurate ACF estimates: TWI reduces the biases in Kalman smoothing when $h=0$ and in linear interpolation when $h=1$, and $h=2$. For the nonlinear Model 3 (TAR), the ACFs estimated from the TWI-imputed series exhibit the lowest RMSEs, with substantial improvements over the initializations. Notably, the scalar filter also yields accurate ACFs for $h=1,2$. Figure S1 plots the histogram of the estimated ACFs at lag 1 (that is, $\gamma(\hat{\mathbf{w}}^{(i)},1)$), which shows how TWI effectively nullifies the biases caused by the initializations.
The advantage of TWI becomes even more pronounced in nonstationary models 4 (I(1)) and 5 (CYC), for which it achieves low ACF estimation errors. The scalar filter performs poorly for these models, which is likely the consequence of using many indicators in the intervention analysis. Among the benchmarks, Kalman filter produces the best result for these two models, but it can be further improved by the proposed TWI procedure.
@@
Figure S1: Histogram of estimated autocovariance function at lag 1 when the data are generated from Model 3 (TAR). The blue line is the smoothed density estimated from the histogram of the ACFs estimated from the ground truth.
@@
Model $h$ Linear Spline Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
@@
S3.3 Autoregressive modeling with groundwater data
As our second downstream task for the groundwater data analysis, we fit an AR(6) model using data imputed by the methods considered. The training data consists of the early period between October, 1992 and September, 2002, amounting to 120 data points per series, including imputed values. Then, we assess the impact of imputation quality by the out-of-sample prediction performance in predicting the remaining data. The missing data in the test data are removed for fair comparison. Figure S3 presents, in violin plots, the distribution of the logarithm of the root mean squared prediction errors ($\log$ RMSE) for each method across the 176 series under study. The TWI-type methods effectively alleviate some large RMSEs caused by the benchmark imputation methods and generally produce lower RMSEs. This shows the series imputed by TWI have marginal distributions that are stable (across time), meaning the method does not artificially induce structural shifts. This further corroborates the practical utility of TWI.
Figure S3: The distributions of the logarithm of the root mean squared prediction errors (RMSEs) for each method across the 176 series, ranked by median from left to right.

### v3

@@ 7 Conclusion
The file contains the proofs for the results presented in Section 4 and some technical results. It also documents the scalar filter algorithm of Peña and Tsay, (2021) and complementary numerical results, including the simulation results for autocovariance estimation and for a compositional time series DGP. Finally, it includes the implementation details of the groundwater data analysis.
Supplementary material to “Temporal Wasserstein Imputation: A Versatile Method for Time Series Imputation”
Shuo-Chieh Huang, Tengyuan Liang, and Ruey S. Tsay
This supplementary material contains three sections. Section S1 collects the proofs of the propositions presented in main text as well as some technical results. Section S2 documents the scalar filter algorithm of Peña and Tsay, (2021) for completeness. Section S3 contains complementary numerical results, including the simulation results for autocovariance estimation and imputing data generated from a compositional time series model. It also includes the details of implementation for the methods employed in the groundwater data analysis.
@@
Since $\ell$ is convex, $\ell^{\prime\prime}\geq 0$. From (S1.1)–(S1.3), it is readily seen that the Hessian matrix $\mathbf{H}=\partial^{2}F_{0}/\partial\mathbf{w}^{2}$ is weakly diagonally dominant with nonnegative diagonal entries. It follows from the Gershgorin circle theorem that $\mathbf{H}$ is positive semi-definite. Thus $F_{0}$ is convex in $\mathbf{w}$. ∎
In order to prove Proposition 4, we need the following lemma.
@@
[…]
@@
[…]
@@
In the rest of this section, the derivation of (13) in Section 4.2 is provided.
Proof of (13) in Section 4.2.
From (9)–(11) in the main text, we can compute the marginal distributions implied by the imputed series, given as follows. If $t<n_{1}$,
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,1))=$ $(1-\frac{1}{k_{1}})\lambda_{2}(1-q)+\frac{1}{k_{1}}\lambda_{2}a_{1},$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(1,0))=$ $(1-\frac{1}{k_{1}})\lambda_{1}p+\frac{1}{k_{1}}\lambda_{1}b_{1},$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,1))=$ $(1-\frac{1}{k_{1}})\lambda_{2}q+\frac{1}{k_{1}}\lambda_{2}(1-a_{1}),$
$\mathbb{P}((\hat{w}_{t},\hat{w}_{t-1})=(0,0))=$ $(1-\frac{1}{k_{1}})\lambda_{1}(1-p)+\frac{1}{k_{1}}\lambda_{1}(1-b_{1}),$
and for $t\geq n_{1}$, the marginal distribution can be obtained by replacing $a_{1},b_{1},k_{1}$ by $a_{2},b_{2},k_{2}$ in the above. Since TWI equates the marginal distributions before and after $n_{1}$, we have
$(1-\frac{1}{k_{1}})\lambda_{2}(1-q)+\frac{1}{k_{1}}\lambda_{2}a_{1}=$ $(1-\frac{1}{k_{2}})\lambda_{2}(1-q)+\frac{1}{k_{2}}\lambda_{2}a_{2},$
$(1-\frac{1}{k_{1}})\lambda_{1}p+\frac{1}{k_{1}}\lambda_{1}b_{1}=$ $(1-\frac{1}{k_{2}})\lambda_{1}p+\frac{1}{k_{2}}\lambda_{1}b_{2},$
$(1-\frac{1}{k_{1}})\lambda_{2}q+\frac{1}{k_{1}}\lambda_{2}(1-a_{1})=$ $(1-\frac{1}{k_{2}})\lambda_{2}q+\frac{1}{k_{2}}\lambda_{2}(1-a_{2}),$
$(1-\frac{1}{k_{1}})\lambda_{1}(1-p)+\frac{1}{k_{1}}\lambda_{1}(1-b_{1})=$ $(1-\frac{1}{k_{2}})\lambda_{1}(1-p)+\frac{1}{k_{2}}\lambda_{1}(1-b_{2}).$
After some algebraic manipulations, these equations simplify to (13) in the main text. ∎
@@
where $\hat{\mathbf{w}}^{(i)}=(\hat{w}_{0}^{(i)},\ldots,\hat{w}_{n-1}^{(i)})$ and $\gamma_{\mathbf{x}}(h)=\frac{1}{1000}\sum_{i=1}^{1000}\gamma(\mathbf{x}^{(i)},h)$ with $\mathbf{x}^{(i)}=(\tilde{x}_{0}^{(i)},\ldots,\tilde{x}_{n-1}^{(i)})$.
Table S1 reports $\mathrm{RMSE}_{\gamma}(h)$ for $h=0,1,2$. For Model 1 (AR), most methods perform reasonably well. Linear interpolation produces accurate ACF estimates since the positive AR coefficient aligns well with linearity, while the scalar filter exhibits largest estimation errors. For Model 2 (ARMA), TWI improves from linear and Kalman initializations and yields more accurate ACF estimates: TWI reduces the biases in Kalman smoothing when $h=0$ and in linear interpolation when $h=1$, and $h=2$. For the nonlinear Model 3 (TAR), the ACFs estimated from the TWI-imputed series have RMSEs that are among the lowest, with substantial improvements over the initializations. Notably, the scalar filter also yields accurate ACFs for $h=1,2$. Figure S1 plots the histogram of the estimated ACFs at lag 1 (that is, $\gamma(\hat{\mathbf{w}}^{(i)},1)$), which shows how TWI effectively nullifies the biases caused by the initializations.
For DGP 4 (I(1)) and DGP 5 (CYC), TWI again offers great improvement over initializations. Noe that for DGP 5 (CYC), which generates a structural time series with deterministic cycles, iSSA is very effective at least under missing pattern I. However, perhaps due to the small window length caused by the high missing ratio, the performance of iSSA under missing pattern II is severely compromised (see also Figure S2, which shows TWI captures the underlying dynamic well). The scalar filter performs poorly for these models, which is likely the consequence of using many indicators in the intervention analysis. Among the benchmarks, Kalman smoothing produces the best result for these two models, which can benefit from further improvement provided by the proposed TWI procedure.
@@
Figure S1: Histogram of estimated autocovariance function at lag 1 when the data are generated from Model 3 (TAR). The blue line is the smoothed density of the histogram of the ACFs estimated from the ground truth.
(a) $h=1$
(b) $h=2$
(c) $h=6$
(d) $h=10$
Figure S2: Scatter plots of $(\tilde{x}_{t-h},\tilde{x}_{t})$ of the original data and $(\hat{w}_{t-h},\hat{w}_{t})$ of the imputed series, when the data are generated from Model 5 (CYC) under missing pattern $I$. $h\in\{1,2,6,10\}$.
@@
Model $h$ Linear iSSA Kalman ScalarF TWIlin $k$-TWIlin TWIKal $k$-TWIKal
@@
S3.3 Implementation details in groundwater data analysis
The methods employed in the groundwater data analysis are detailed as follows. First, linear interpolation, mean imputation, and Kalman smoothing are implemented using the imputeTS package (Moritz and Bartz-Beielstein,, 2017) in R. For the iterative singular spectrum analysis (iSSA) method, we implement it via the Rssa package (Golyandina et al.,, 2018). There are three parameters to tune for the iSSA: the initialization, the window length, and the rank in the reconstruction step. Mean imputation is adopted here as the initialization, which is a common choice in practice. We set the window length to 6, which is roughly the maximum window length allowed when the missing ratio is high. Finally, to determine the rank in the reconstruction step, we employ a cross-validation approach. Specifically, we randomly mask 10% of the observed data. Then, for each rank $r\in\{1,2,\ldots,6\}$, we apply the iSSA for imputation and select the rank $r^{*}$ that yields the most accurate imputations on the masked data, measured by the mean squared error. Finally, the iSSA is implemented again with the full sample and the selected rank.
For the TWI-type methods, as discussed in the main text, we first apply kernel smoothing to estimate a trend function $f(t)$. We use the locfit package (Loader,, 2024) with the default tricube kernel and the bandwidth is set to 12 months. This step can be implemented regardless of the missing pattern. Then the ($k$-)TWI is applied to the residuals $r_{t}=x_{t}-f(t)$, which has the same missing pattern as $\{x_{t}\}$, to obtain $\{\hat{r}_{t}\}$. For the TWI, the cut-off point $n_{1}$ is set to $\lfloor 0.4n\rfloor$ as in the simulation section. For $k$-TWI, the cut-offs are $\{0.3n,0.7n,0.5n\}$. The simple last observation carried forward (LOCF) is used as initialization for both TWI and $k$-TWI. Finally, the imputation is $\hat{x}_{t}=\hat{r}_{t}+f(t)$.